In a constant-pressure calorimeter, 50.0 mL of 0.310 M Ba(OH)2 was added to 50.0 mL of 0.620 M HCI The reaction caused the temperature of the solution to rise from 23.56 ? to 27.78 ?, what is ?? for this reaction (KJ released per mole of H2O produced)? Assume that the total volume is the sum of the individual volumes and that the density and heat capacity are equal to that of water Number kJ/ mol H,O
Answer #1

no of moles of Ba(OH)2 = molarity * volume in L

                                       = 0.31*0.05   = 0.0155moles

no of moles of HCl          = molarity * volume in L

                                       = 0.62*0.05 = 0.031 moles

Ba(OH)2(aq) + 2HCl(aq) ------------> BaCl2(aq) + 2H2O(l)

Total voume of solution = 50 + 50 = 100ml

The density of solution = 1g/ml

mass of solution = volume of solution * density of solution

                            = 100*1 = 100g

q = mc\DeltaT

      = 100*4.184*(27.78-23.56)   = 1765.65J    = 1.76565KJ

\DeltaH   = -1.76565KJ

the energy converted to Kj/mole

\DeltaH   = -1.76565/0.031   =- 57KJ/mole >>>>answer